To the nearest hundreth, find the circumference of a circle with a radius of 7 centimeters.
I keep getting 43.96 and am told that it is wrong.

Answer by scott8148(6628)

c = 2 * pi * r = 2 * 3.14159265... * 7 = 43.98229715...

not enough decimal places for pi or rounding error?